# Interactive Agents to Overcome Ambiguity in Software Engineering

### v1

@@ Abstract
AI agents are increasingly being deployed to automate tasks, often based on ambiguous and underspecified user instructions. Making unwarranted assumptions and failing to ask clarifying questions can lead to suboptimal outcomes, safety risks due to tool misuse, and wasted computational resources. In this work, we study the ability of LLM agents to handle ambiguous instructions in interactive code generation settings by evaluating proprietary and open-weight models on their performance across three key steps: (a) leveraging interactivity to improve performance in ambiguous scenarios, (b) detecting ambiguity, and (c) asking targeted questions. Our findings reveal that models struggle to distinguish between well-specified and underspecified instructions. However, when models interact for underspecified inputs, they effectively obtain vital information from the user, leading to significant improvements in performance and underscoring the value of effective interaction. Our study highlights critical gaps in how current state-of-the-art models handle ambiguity in complex software engineering tasks and structures the evaluation into distinct steps to enable targeted improvements. [^1] Code and data can be accessed at https://github.com/sani903/InteractiveSWEAgents
Software Engineering Agents,Human-AI Interaction,Intent Alignment,Ambiguity Resolution
[…]
gnmagenta
Figure 1: Interactive agents mitigate resource wastage and reduce misalignment in ambiguous settings.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly used as chatbots in task-oriented workflows to improve productivity (Peng et al., 2023; Brynjolfsson et al., 2023), with the user providing a task instruction which the model completes. Due to the interactive nature of chatbots, the performance depends on the information provided in the user’s prompt. Users often provide non-descriptive instructions, which poses critical challenges in successfully completing the task (Chowdhury et al., 2024). The ambiguity can lead not only to erroneous outcomes, but also to significant safety issues (Kim et al., 2024; Karli & Fitzgerald, 2023).
This ambiguity can lead to more severe consequences in task automation scenarios, where AI agents are equipped with powerful tools (Wang et al., 2024b; Lu et al., 2024; Huang et al., 2024; Zhou et al., 2024a). In software engineering settings, agents must navigate complex codebases, make architectural decisions, and modify critical systems—all while operating with potentially incomplete or ambiguous instructions. When human developers face such ambiguity, they engage in clarifying dialogue to gather missing context (Testoni & Fernández, 2024; Purver, 2004). However, current AI systems often proceed with incomplete understanding, leading to costly mistakes and misaligned solutions, as demonstrated in Figure 1.
In this work, we systematically evaluate the interaction capabilities of commonly used open and proprietary LLMs when addressing underspecified instructions in agentic code settings (§2). We examine three research questions to address the problem for code generation.
@@
    Interactive problem solving: Can LLMs appropriately leverage interaction with the user to improve performance in ambiguous settings?
@@
    Detection of ambiguity: Can LLMs identify whether a given task description is underspecified and ask clarifying questions?
@@
    Question quality: Can LLMs generate meaningful and targeted questions that gather the necessary information to complete the task?
Figure 2: The three settings in order: Full, Hidden, and Interaction
We evaluate the research questions separately to ensure independence between them. We use the Github issues from SWE-Bench Verified (Chowdhury et al., 2024) to simulate well-specified inputs, and the summarized variants of the same Github issues as underspecified inputs for the experiments. A simulated user (Xu et al., 2024; Zhou et al., 2024b), equipped with the full, well-specified issue, simulates real conversations where the user has additional context, which is provided only when prompted with the appropriate questions. This multi-stage approach allows for targeted improvements in individual aspects, offering a pathway to enhance overall system performance.
Through our evaluations across the different settings, we find that interactivity can boost performance on underspecified inputs by up to 74% over the non-interactive settings but the performance varies between models (§3). LLMs default to non-interactive behavior without explicit encouragement, and even with it, they struggle to distinguish between underspecified and well-specified inputs. Claude Sonnet 3.5 is the only evaluated LLM that achieves notable accuracy (84%) in making this distinction. Prompt engineering offers limited improvement, and its effectiveness varies across models (§4). When interacting, LLMs generally pose questions capable of extracting relevant details, but some models, such as Llama 3.1 70B, fail to obtain sufficient specificity (§4). In summary, this study underscores the importance of interactivity in LLMs for agentic workflows, particularly in real-world tasks where prompt quality varies significantly.
@@ 2.1 Dataset
Given that the Verified subset contains only sufficiently specified issues, we assume that these issues do not require disambiguation. Therefore, for each SWE-Bench Verified issue, we consider two forms, as shown in Figure 2:
@@
    Underspecified issue: A summarized version generated using GPT-4o, where the model is asked to preserve specific terminology is preserved but reduce the amount of detailed content (complete prompt in §A.1.3).
@@ 2.2 Agentic Framework
We use Claude Sonnet 3.5 (Anthropic, 2024b) as one of the proprietary models due to its superior performance on SWE-Bench. Claude Haiku 3.5 (Anthropic, 2024a) is included as the second proprietary model to investigate the impact of model parameterization, as both models likely share similar training methodologies but differ significantly in the number of parameters. Additionally, we evaluate Llama 3.1 70B-Instruct (Llama team, 2024) and Deepseek-v2 (DeepSeek-AI, 2024) as two open-weight frontier models.
@@
We employ GPT-4o (Ahmad & OpenAI, 2024) as a user proxy to simulate user-agent interactions (Xu et al., 2024; Zhou et al., 2024a). The user proxy is provided with the fully specified version of the task, allowing the coding agent to extract the necessary information through interaction. It is instructed to respond based solely on the information available in the full issue and will reply with I don’t have that information if relevant details are missing. This approach prevents the user proxy from hallucinating incorrect information and encourages clear, negative responses when needed. The full prompt is shown in §A.1.2.
@@ 2.3 Study Design
  • Full Setting: This is the traditional SWE-Bench setting for resolving GitHub issues. The coding agent is provided with the fully specified task and the interaction is disabled. It represents the agent’s performance in an unambiguous scenario, where the agent has access to full information, simulating ideal conditions.
  • Hidden setting: A summarized version of the issue is provided to the coding agent with the user-agent interaction disabled to mimic the lack of detail that can occur in task descriptions. We do not give any interaction-related instructions, and all models default to non-interactive behavior. Specific details are hidden from the coding agent.
  • Interaction Setting: The coding agent receives a summarized task, while the user proxy model gets the fully specified task. Interaction is enabled through prompting, allowing the agent to query the proxy for specific details. The models do not interact with the user without an explicit prompt. In addition to the full issue, the proxy has access to file locations that need modification and can provide them when queried. This setup allows us to evaluate which models proactively seek navigational information and examine how this interaction influences the success of the solution process across models.
@@ 3 RQ1: Interactive Problem Solving
Effectively addressing ambiguity requires a model to integrate information from user interactions to form a clear plan and successfully solve the task. Our first experiment holistically evaluates the model’s ability to leverage interaction and improve performance. The model must not only process the initial task description, but also query users to extract relevant details while filtering out irrelevant information.
@@ 3.1 Experimental Setup
The hypothesis of the experiment is that different language models will exhibit varying performance with interaction based on their incorporation of the provided information, leading to different levels of improvement over the Hidden setting. We evaluate the models across the three settings and conduct two Wilcoxon-Signed Rank tests with a significance level of 0.05 to determine significant performance differences between the Hidden and Interaction settings, and between the Interaction and Full settings for every model. Here, we modify the prompt to make interaction with the user compulsory in the Interaction setting [^2] Without compulsory interaction, the model defaults to non-interactive behavior for most issues, as seen in the Hidden setting. Full prompt in §A.1.2. Ideally, the Interaction setting should approach the performance of the full setting. The coding agent has a maximum of 30 turns to generate a solution patch.
3.2 Leveraging Interaction in Ambiguity
In this experiment, each model is tested in the Hidden, Interaction, and Full settings to evaluate its ability to leverage interaction and optimize performance on underspecified issues. The results, as shown in Figure 3, confirm the expected increase in resolve rates as more information becomes available to the agent. While the difference between the Hidden and Interaction settings is significant for every model (Table 1), emphasizing the impact of interaction on the trajectory, the performance gap between the Interaction and Full settings is also significant across all models, highlighting the unrealized potential. Specifically, for the Hidden vs. Interaction settings, proprietary models show stronger evidence of a significant difference. These results suggest that the ability to leverage interaction varies across models, with proprietary models demonstrating greater effectiveness in utilizing interaction compared to open-weight models.
@@ 3.2 Leveraging Interaction in Ambiguity
Figure 3: Resolve rates (in %) across different settings: Hidden (underspecified issues), Interaction (underspecified issues with user interaction), and Full (fully specified issues)

### v2

# Ambig-SWE: Interactive Agents to Overcome Underspecificity in Software Engineering

@@ Abstract
AI agents are increasingly being deployed to automate tasks, often based on underspecified user instructions. Making unwarranted assumptions to compensate for the missing information and failing to ask clarifying questions can lead to suboptimal outcomes, safety risks due to tool misuse, and wasted computational resources. In this work, we study the ability of LLM agents to handle underspecified instructions in interactive code generation settings by evaluating proprietary and open-weight models on their performance across three key steps: (a) detecting underspecificity, (b) asking targeted clarification questions, and (c) leveraging the interaction to improve performance in underspecified scenarios. We introduce Ambig-SWE, an underspecified variant of SWE-Bench Verified, specifically designed to evaluate agent behavior under ambiguity and interaction. Our findings reveal that models struggle to distinguish between well-specified and underspecified instructions. However, when models interact for underspecified inputs, they effectively obtain vital information from the user leading to significant improvements in performance, up to 74% over the non-interactive settings, underscoring the value of effective interaction. Our study highlights critical gaps in how current state-of-the-art models handle missing information in complex software engineering tasks and structures the evaluation into distinct steps to enable targeted improvements [^1] Code and data can be accessed at https://github.com/sani903/InteractiveSWEAgents.
@@ 1 Introduction
Figure 1: Interactive agents reduce resource wastage and misalignment in underspecified settings.
Large Language Models (LLMs) are increasingly used as chatbots in task-oriented workflows to improve productivity (Peng et al., 2023; Brynjolfsson et al., 2023), with the user providing a task instruction which the model completes. Due to the interactive nature of chatbots, the performance depends on the information provided in the user’s prompt. Users often provide non-descriptive instructions, which poses critical challenges in successfully completing the task (Chowdhury et al., 2024). The missing information can lead not only to erroneous outcomes, often based on hallucinations, but also to significant safety issues (Kim et al., 2024; Karli and Fitzgerald, 2023).
This underspecificity can lead to severe consequences in task automation, where AI agents are equipped with powerful tools (Wang et al., 2024b; Lu et al., 2024; Huang et al., 2024; Zhou et al., 2024a). In software engineering settings, agents navigate complex codebases, make architectural decisions, and modify critical systems—all while operating with potentially incomplete instructions. When human developers face such lack of information, they engage in clarifying dialogue to gather context (Testoni and Fernández, 2024; Purver, 2004). However, current AI systems proceed with incomplete understanding, leading to costly mistakes and misaligned solutions.
In this work, we systematically evaluate the interaction capabilities of commonly used open and proprietary LLMs when addressing underspecified instructions in agentic code settings (§2). We define underspecificity as missing information that would prevent an expert from being able to create a successful solution, using the same definition as SWE-Bench Verified annotation rubric. Previous work on underspecificity (Chen et al., 2025; Kim et al., 2024) typically focuses on cases where only a single detail is missing. In contrast, real-world agentic tasks often involve multiple, interdependent gaps in specification that emerge over the course of a trajectory—spanning file locations, design decisions, and constraints—making the problem substantially harder and motivating new evaluation frameworks. Our work makes the following contributions:
@@
    Evaluating underspecificity in complex agentic tasks. We extend SWE-Bench Verified with underspecified variants of GitHub issues and introduce an interactive evaluation framework, Ambig-SWE, where agents can query a simulated user (Xu et al., 2024; Zhou et al., 2024b) holding the full specification. This design enables controlled study of how agents handle different forms and levels of underspecificity in realistic multi-step workflows. We also compare against the standard SWE-Bench setting and a non-interactive underspecified setting to analyze differences in agent trajectories.
@@
    Analysis of interaction capabilities We break down resolution under underspecificity into three fundamental capacities: (i) detecting when instructions are incomplete, (ii) acquiring the missing details through targeted clarification, and (iii) leveraging the interaction to successfully complete the task. We design evaluations for each capacity and measure performance across proprietary and open-weight models.
@@
    Empirical insights for agent design Our experiments show that interactivity can recover performance lost to underspecificity, but most LLMs default to non-interactive behavior and struggle with robust detection. We identify actionable clarifying questions as the main driver of performance gains, providing concrete guidance for future model and agent design.
The multi-stage evaluation allows for targeted improvements in individual aspects, offering a pathway to enhance overall system performance. Through our evaluations across the different settings, we find that interactivity can boost performance on underspecified inputs by up to 74% over the non-interactive settings, though performance varies across models (§3). LLMs default to non-interactive behavior without explicit encouragement, and even with it, they struggle to distinguish between underspecified and well-specified inputs. Claude Sonnet 4 and Claude Sonnet 3.5 are the only evaluated LLMs that achieve notable accuracy (89% and 84%, respectively) in making this distinction. Prompt engineering offers limited improvement, and its effectiveness varies across models (§4). When interacting, LLMs generally pose questions capable of extracting relevant details, but some models, such as Llama 3.1 70B, fail to obtain sufficient specificity (§4). As models grow more capable (e.g., from Claude Sonnet 3.5 to Claude Sonnet 4), interaction provides diminishing returns, suggesting current training practices may not adequately leverage clarification. In summary, this study underscores the importance of interactivity in LLMs for agentic workflows, particularly in real-world tasks where prompt quality varies significantly.
@@ 2.1 Dataset
Given that the Verified subset contains only sufficiently specified issues as per human annotations, we assume that these issues do not require more information. Therefore, for each SWE-Bench Verified issue, we consider two forms, as shown in Figure 2:
@@
    Underspecified issue: A synthetic version generated using GPT-4o, where the model is asked to preserve specific terminology is preserved but reduce the amount of detailed content (complete prompt in Appendix §A.2.3).
We conduct an analysis comparing annotated underspecified issues in SWE-Bench with our generated underspecified issues using distributional difference analysis (Zhong et al., 2023) to identify how the underspecification in our generations varies from real user issues. Our findings show that natural underspecified issues have more (1) concrete technical details (code snippets, error messages, file/line references), (2) reproducibility information, (3) links to external references, and (4) conversational fragments (stream of thought, incomplete sentences)
In contrast, our generated issues did not have any particular additional features—they do not have traits that are statistically more common than natural issues. Our approach uses more aggressive information removal, specifically targeting code snippets and error messages. However, there are naturally occurring underspecified issues that are similarly vague as well. The other differences (external links, conversational style) may not directly impact agent performance since agents cannot access external information.
To assess the extent of information loss in the underspecified issues of our dataset, we provide quantitative metrics in the Appendix §A.2.3. For a concrete specification of missing information between the fully specified and the underspecified issue, we use an LLM to annotate the differences [^2] LLM annotations for underspecification are provided in the supplementary materials.. A qualitative evaluation of the summaries confirms the findings of the distributional difference analysis. We did not evaluate on naturally underspecified SWE-Bench examples because they lack the paired ground truth (complete specifications) necessary for causal measurement of interaction impact. Without verified correct specifications, we cannot determine whether performance improvements result from resolving genuine underspecification versus other confounding factors.
Figure 2: The three settings in order: Full, Hidden, and Interaction.
@@ 2.2 Agentic framework
We evaluate two proprietary models from the same family—Claude Sonnet 3.5 and its successor Claude Sonnet 4 (Anthropic, 2024b; PBC, 2025)—to study how improvements in model capabilities influence interaction behavior. We also include Claude Haiku 3.5 (Anthropic, 2024a), which shares similar training methodology but differs substantially in parameter count, allowing us to examine the effect of model scale.
For open-weight models, we evaluate Llama 3.1 70B-Instruct (Llama team, 2024), Deepseek-v2 (DeepSeek-AI, 2024), and Qwen 3 Coder 480B. Qwen 3 Coder achieves performance comparable to Claude Sonnet 4 on SWE-Bench, enabling a comparison of interaction patterns between models with similar task-solving capabilities.
@@
Following related works which used LLMs to simulate users with full information (Li et al., 2024), we employ GPT-4o (Ahmad and OpenAI, 2024) as a user proxy to simulate user-agent interactions. This design choice is informed by prior work showing that LLMs can approximate simple user behaviors and produce natural-sounding responses in controlled settings (Xu et al., 2024; Zhou et al., 2024a). The goal is not to simulate real users but provide the information injection to the trajectory and analyze model behaviors. The proxy receives the full issue and responds only using information explicitly present in it, preserving the original knowledge boundaries of the issue reporter. If a queried detail is missing, the proxy responds with I don’t have that information, thereby avoiding hallucinations. This conservative design makes it possible to isolate the agent’s ability to detect and recover from missing information. The full prompt is provided in §A.2.2.
@@ 2.3 Study design
  • Full: This is the traditional SWE-Bench setting. The coding agent is provided with the fully specified task and the interaction is disabled. It represents the agent’s performance in an ideal scenario, where the agent has access to full information.
  • Hidden: A summarized version of the issue is provided to the coding agent with the user-agent interaction disabled to mimic the lack of detail that can occur in task descriptions. We do not give any interaction-related instructions, and all models default to non-interactive behavior. Specific details are hidden from the coding agent.
  • Interaction: The coding agent receives a summarized task, while the user proxy model receives the fully specified task. Interaction is enabled through prompting, allowing the agent to query the proxy for specific details. The models do not interact without an explicit prompt. In addition to the full issue, the proxy has access to file locations that need modification and can provide them when queried. This setup allows us to evaluate which models proactively seek navigational information and examine how this interaction influences the success of the solution process across models.
@@ 3 RQ1: Interactive problem solving
Can LLMs appropriately leverage interaction with the user to improve performance in underspecified settings? Effectively addressing missing information requires a model to integrate information from user interactions to form a clear plan and successfully solve the task. Our first experiment holistically evaluates the model’s ability to leverage interaction and improve performance. The model must not only process the initial task description, but also query users to extract relevant details while filtering out irrelevant information.
@@ 3.1 Experimental setup
The hypothesis of the experiment is that different language models will exhibit varying performance with interaction based on their incorporation of the provided information, leading to different levels of improvement over the Hidden setting. We evaluate the models across the three settings and conduct two Wilcoxon-Signed Rank tests (Appendix §A.3.1) with a significance level of 0.05 to determine significant performance differences between the Hidden and Interaction settings, and between the Interaction and Full settings for every model. Here, we modify the prompt to make interaction with the user compulsory in the Interaction setting [^3] Without compulsory interaction, the model defaults to non-interactive behavior for most issues, as seen in the Hidden setting. Full prompt in §A.2.2. Ideally, the Interaction setting should approach the performance of the full setting.
@@
Figure 3: Resolve rates (in %) across different settings: Hidden (underspecified issues), Interaction (underspecified issues with user interaction), and Full (fully specified issues).
By default, coding agents are restricted to 30 interaction turns to produce a solution patch; however, Claude Sonnet 4 and Qwen 3 Coder are allocated up to 100 turns to account for their greater reasoning and planning capacity. In this experiment, each model is tested in the Hidden [^4] Claude Sonnet 4 is evaluated on a subset of 100/500 instances in the Hidden setting. The model compensates for the lack of information with increased exploration and solution attempts leading to substantially higher evaluation costs. The findings are still statistically significant (Table 4)., Interaction, and Full settings to evaluate its ability to leverage interaction and optimize performance on underspecified issues. The results, as shown in Figure 3, confirm the expected increase in resolve rates for all models as more information becomes available to the agent. The difference between the Hidden and Interaction settings is significant for all evaluated models (Table 4), emphasizing the impact of interaction on task completion. The performance gap between the Interaction and Full settings is also significant across all models, highlighting the unrealized potential. Specifically, for the Hidden vs. Interaction settings, proprietary models show stronger evidence of a significant difference. These results suggest that the ability to leverage interaction varies across models, with proprietary models generally demonstrating greater effectiveness in utilizing interaction compared to open-weight models. However, as open-weight models improve, they can even outperform proprietary models with interaction, as demonstrated by Qwen 3 Coder.
3.2 Leveraging interaction in underspecificity